Write the slope-intercept form of the equation of the line described.
through (5,-5), parallel to y=-8/5x+2.

Answer by TimothyLamb(4379)

---
given:
(5,-5)
parallel to y=-8/5x+2
---
point = (5,-5)
slope = -8/5 = -1.6
---
to solve the linear equation, copy and paste this (the given point and the parallel slope):
5,-5,-1.6
into the "Point-Slope form: x1 y1 m" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (1.875, 0)
y-intercept = (0, 3)
slope = -1.6
---
answer:
the slope-intercept form of our new line:
y = -1.6x + 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php